Find the polar equations of the lines or curves whose Cartesian equations are
(a) (x x a)
2 + (y y a)
2 = a
2
(b) x
2/a2 + y
2/b2 = 1 (c) y = 2x (d) y = x
2
(e) xy = 4 (f) y = 4 (g) 3x x y + 2 = 0.
1
Expert's answer
2020-04-24T18:20:46-0400

Note: for verification I will use a graphical online calculator -https://www.desmos.com/calculator

As we know the Cartesian coordinates are related to polar related relationships



"\left\{\begin{array}{l}
x=r\cdot\cos\theta\\
y=r\cdot\sin\theta
\end{array}\right."

( More information: https://en.wikipedia.org/wiki/Polar_coordinate_system )

This is enough to solve this task.


(с) "y=2x"

Then,



"r\cdot\sin\theta=2r\cdot\cos\theta\to
\left[\begin{array}{l}
r=0\\
\left.\sin\theta=2\cos\theta\right|\div\left(\cos\theta\right)
\end{array}\right.\\[0.3cm]
\left[\begin{array}{l}
r=0-\text{describes the origin}\\
\tan\theta=2-\text{describes other points of the line}
\end{array}\right."

(d) "y=x^2"



"r\cdot\sin\theta=r^2\cdot\cos^2\theta\to
\left[\begin{array}{l}
r=0\\
\left.\sin\theta=r\cos^2\theta\right|\div\left(\cos^2\theta\right)
\end{array}\right.\\[0.3cm]
\left[\begin{array}{l}
r=0-\text{describes the origin}\\[0.3cm]
r=\displaystyle\frac{\sin\theta}{\cos^2\theta}-\text{describes other points of parabola}
\end{array}\right.\\[0.3cm]"




(e) "xy=4"



"\left.\left(r\cdot\sin\theta\right)\cdot\left(r\cdot\cos\theta\right)=4\right|\times(2)\longrightarrow\\[0.3cm]
r^2\cdot\left(\underbrace{2\sin\theta\cdot\cos\theta}_{sin2\theta}\right)=8\longrightarrow\\[0.3cm]
r^2=\frac{8}{\sin2\theta}\longrightarrow\boxed{r=\sqrt{\frac{8}{\sin2\theta}}}"





(f) "y=4"



"r\cdot\sin\theta=4"




(g) "3x-y+2=0"



"3r\cdot\cos\theta-r\cdot\sin\theta+2=0\longrightarrow\\[0.3cm]
r\cdot\left(3\cos\theta-\sin\theta\right)=-2\longrightarrow\\[0.3cm]
r=\frac{-2}{3\cos\theta-\sin\theta}"





(b) "x^2\/a^2+y^2\/b^2=1"



"\left.\frac{\left(r\cdot\cos\theta\right)^2}{a^2}+\frac{\left(r\cdot\sin\theta\right)^2}{b^2}=1\right|\times\left(a^2b^2\right)\\[0.3cm]
r^2\cdot\left(b^2\cdot\cos^2\theta+a^2\cdot\sin^2\theta\right)=a^2b^2\\[0.3cm]
r=\sqrt{\frac{a^2b^2}{b^2\cdot\cos^2\theta+a^2\cdot\sin^2\theta}}\\[0.3cm]
r=\frac{ab}{\sqrt{b^2\cdot\cos^2\theta+a^2\cdot\sin^2\theta}}"

Note: for example, I took "a=2" and "b=3"




(a) "(x-a)^2+(y-a)^2=a^2"



"\left(r\cdot\cos\theta-a\right)^2+\left(r\cdot\sin\theta-a\right)^2=a^2\\[0.3cm]
r^2\cdot\left(\underbrace{\cos^2\theta+\sin^2\theta}_{=1}\right)-2ar\cdot\left(\cos\theta+\sin\theta\right)+2a^2=a^2\\[0.3cm]
r^2-2ar\sqrt{2}\cdot\left(\frac{\sqrt{2}}{2}\cos\theta+\frac{\sqrt{2}}{2}\sin\theta\right)+a^2=0\\[0.3cm]
r^2-2\sqrt{2}ar\cdot\cos\left(\theta-\frac{\pi}{4}\right)+a^2=0"

Note: for example, I took "a=2"